OpenStudy (anonymous):

A local pizzeria offers 11 toppings for their pizzas and you can choose any 5 of them for one fixed price. How many different types of pizzas can you order with 5 toppings? A. 462 B. 120 C. 55,440 D. 332,640

OpenStudy (zarkon):

\[{11\choose 5}\]

OpenStudy (anonymous):

is it A B C or D

OpenStudy (anonymous):

\[{11\choose 5}=\frac{11\times 10\times 9\times 8\times 7}{5\times 4\times 3\times 2}\]cancel first, multiply last

OpenStudy (anonymous):

\[{11\choose 5}=\frac{11\times 10\times 9\times 8\times 7}{5\times 4\times 3\times 2}=11\times 3\times 2\times 7\] is the method. if you only want the answer, use a calculator http://www.wolframalpha.com/input/?i=11+choose+5

OpenStudy (anonymous):

thank you!!